OpenStudy (anonymous):

Solve the system of equations. 3x + 4y =8 -2x + 5y = 3

OpenStudy (migitmack):

Are these answers together or seperate

OpenStudy (migitmack):

The answer for the first one is 8/3-4y/3

OpenStudy (migitmack):

The answer for the second one is 5y/2-3/2

OpenStudy (anonymous):

well its 2 equation but in one problem..im suppose to solve the system of equations

OpenStudy (anonymous):

can you guys set it up?

OpenStudy (migitmack):

/ This sign means fraction sign

OpenStudy (anonymous):

ok..im saying i cant just put a answer..can you show me how you solved it??

OpenStudy (anonymous):

yesterday i showed you substitution, now i'll show you elimination: 3x + 4y = 8....(i) -2x +5y = 3....(ii) multiply equation (i) by 5 and equation (ii) by 4 inorder to be able to get same value of y 5 (3x + 4y = 8) = 15x +20y = 40 ......(i') 4 (-2x +5y = 3) = -8x +20y = 12......(ii') subtract (ii) from (i) 15x -(-8x) +20y - 20y = 40-12 23x = 28 x = 28/23 - = 1 5/23 from (i) y = (8-3x)/4 y = 1 2/23 solution (x,y) = (1 5/23, 1 2/23)

OpenStudy (anonymous):

:)) thank youu

OpenStudy (anonymous):

:)